OpenStudy (anonymous):

1 over z - 1/2z - 1/5z = 10 over 21

OpenStudy (anonymous):

\[\frac{1}{z}-\frac{1}{2z}-\frac{1}{5z}=\frac{10}{21}\] Multiplying everything by 10z to get rid of the denominators we obtain: \[\frac{10z}{z}-\frac{10z}{2z}-\frac{10z}{5z}=\frac{100z}{21}\] \[10-5-2 = (\frac{100}{21})z \Rightarrow 3(21) = 100z\] So z is just: \[z=\frac{3(21)}{100} = \frac{63}{100} = .63\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

hmm this isnt right it was supposed to be 3/97

OpenStudy (anonymous):

Typing it directly into wolframalpha.com it gave me the same answer I gave you.